PROBLEMS:

1.The real GDP, and consumption data for Nation B in 1996 is in the following table. The Investment is 25 billions, and the government purchases is 5 billions. Nation B has no international trade.

------------------------------------------------------------------------------------------------------

Real GDP is billions        100            200           300            400          500         600

Consumption in billions 150            230           310            390          470          550

Aggregate Expenditure ______        _____        _____        _____      _____     _____

--------------------------------------------------------------------------------------------------------

a. Complete the AE in the above table.

     The equilibrium expenditure in this nation is ___________ The MPC is______________

     The Multiplier is _____________________________.

2. Fill in the blank spaces in the table according to the classical view.

___________________________________________________________________________________________

State of the   Unemployment rate         SR real GDP                         Wage rate               SRAS   

economy    (high, low, at Un)   (below, above or at natural real GDP) (rise, fall, same)   (rise, fall, same)  

____________________________________________________________________________________________

Recessionary gap        _______           ___________________             __________         ___________            

Inflationary gap          _______           ___________________             __________         ___________            

Long-run equilibrium _______           ___________________               __________         ___________             __________________________________________________________________________________________

3. Complete the table below by computing the average tax rates, given the net tax revenue data in column 2, and 4. And determine which column has the progressive, proportional, or regressive tax system.

---------------------------------------------------------------------------------------------------------

( 1 )           ( 2 )                    ( 3 )             ( 4 )               ( 5 )                ( 6 )

real GDP     net tax revenue   average tax rate    net tax revenue  average tax rate  government spending

---------------------------------------------------------------------------------------------------------

$1000           $100          ____________        $100            ___________        $120

1100          120           ____________          108            ___________         120

1200            145           ____________          115            ___________         120

  1300            175            ____________          120            ___________         120  

1400            210            ____________          123            ___________         120  

---------------------------------------------------------------------------------------------------------

4. Suppose that AD and AS for a hypothetical economy are as shown in the following table. MPC is 0.8.

----------------------------------------------------------------------------------------------------------------------

Amount of real domestic output demanded   Price level    Amount of real domestic output supplied

                              (in billions)                  (price index)                 (in billions)

------------------------------------------------------------------------------------------------------------------------

                                    $300                           $300                           $700

                                      400                            250                             600

                                      500                            200                             500

                                      600                            150                             400

                                      700                            100                             300

-------------------------------------------------------------------------------------------------------------------------

a. In SR, this economy's equilibrium price level is __________ and equilibrium real output is _________

b. Suppose this country's GDP at natural rate of unemployment is $400B.

Is this country in the recessionary gap or in the inflationary gap? ________________________________

c. Suppose Government decided to change its purchases to eliminate the above gap, Government should

__________ spending by ___________ billion. Which curve will government's action change? _______.

Is this an expansionary or contractionary fiscal policy? __________________

Is the resulting equilibrium level of real domestic output the absolute full-capacity real output?________

If not, explain why:____________________________________________________________________
Answer #1

1.

Aggregate Expenditure = Consumption + Investment + Government purchases

Real GDP in billions 100 200 300 400 500 600
Consumption in billions 150 230 310 390 470 550
Aggregate Expenditure in billions 180 260 340 420 500 580

The equilibrium expenditure in this nation is $500 billion. Because here Real GDP = Aggregate Expenditure.

MPC = Change in consumption / Change in Real GDP = 80 / 100 = 0.8

The Multiplier = 1 / (1 - MPC) = 1 / (1 - 0.8) = 5